# Robust $\Lambda$-quantiles and extreme probabilities

### v1

@@ 4.3 Marginal constraints in risk aggregation
The worst-case value of quantiles over the aggregation set has been studied extensively in the literature. It is well known that worst-case value of quantiles in risk aggregation generally does not admit any analytical formula. Typically, the results in the literature either offer explicit expressions for some special type of marginal distributions or provide some upper bounds for the general marginal distributions. We refer to Embrechts et al. (2006) for bounds, and Wang and Wang (2016), Jakobsons et al. (2016) and Blanchet et al. (2020) for the analytical expressions.
Let $\mathcal{M}_{D}$ (resp. $\mathcal{M}_{I}$) denote the set of all distributions with decreasing (resp. increasing) densities. The expressions of the extreme probabilities over uncertainty set $\mathcal{D}_{n}(\mathbf{F})$ are displayed in the following proposition.
@@
For $\mathbf{F}\in\mathcal{M}_{D}^{n}\cup\mathcal{M}_{I}^{n}$, we have
$F_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)=1-H_{\mathbf{F}}(x),~{}F_{\mathcal{D}_{n}(\mathbf{F})}^{+}(x)=H_{\overline{\mathbf{F}}}(-x),~{}x\in\mathbb{R},$ (14)
@@
with $\mathbf{r}=(r_{1},\dots,r_{n})$, $r=\sum_{i=1}^{n}r_{i}$ and $\Theta_{n}=\{(r_{1},\dots,r_{n})\in\mathbb{R}^{n}:\sum_{i=1}^{n}r_{i}<x\},$ and $\overline{\mathbf{F}}=(\overline{F}_{1},\dots,\overline{F}_{n})$ with $\overline{F}_{i}(x)=1-F_{i}(-x),~{}x\in\mathbb{R}$. Moreover, both $\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $F_{\mathcal{D}_{n}(\mathbf{F})}^{+}$ are continuous and attainable.
@@
In light of Theorems 2 and 4 of Blanchet et al. (2020), we have
$\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{+}(F)=H_{\mathbf{F}}^{-1}(1-\alpha),~{}\alpha\in(0,1).$ (15)
Note that $\mathbf{F}\in\mathcal{M}_{D}^{n}\cup\mathcal{M}_{I}^{n}$ implies $F_{1}^{-1},\dots,F_{n}^{-1}$ are continuous over $(0,1)$. Hence, by Lemma 4.5 of Bernard et al. (2014), we have
@@
Moreover, in light of (i) of Corollary 1, we have $\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$ for $\alpha\in(0,1)$. Consequently, $q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})=H_{\mathbf{F}}^{-1}(1-\alpha),~{}\alpha\in(0,1).$ As it is stated in the proof of Theorem 4 of Blanchet et al. (2020), $H_{\mathbf{F}}(x)$ is continuous over $\mathbb{R}$ and strictly decreasing over $(-\infty,\sum_{i=1}^{n}q_{1}^{-1}(F_{i}))$ and is equal to 00 on $[\sum_{i=1}^{n}q_{1}^{-1}(F_{i}),\infty)$. Hence, we have $F_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)=1-H_{\mathbf{F}}(x),~{}x\in\mathbb{R}$.
@@
By (i) of Corollary 1, we have $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{+}(F)=q_{\alpha}^{+}(F_{\mathcal{D}_{n}(\mathbf{F})}^{+})$ for $\alpha\in(0,1)$. Direct computation gives $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{+}(F)=-\sup_{F\in\mathcal{D}_{n}(\overline{\mathbf{F}})}q_{1-\alpha}^{-}(F)$, where $\overline{\mathbf{F}}=(\overline{F}_{1},\dots,\overline{F}_{n})$ with $\overline{F}(x)=\lim_{y\downarrow x}1-F(-x),~{}x\in\mathbb{R}$. Note that $\overline{\mathbf{F}}\in\mathcal{M}_{D}^{n}\cup\mathcal{M}_{I}^{n}$. Hence, using (15) and (16), we have
$q_{\alpha}^{+}(F_{\mathcal{D}_{n}(\mathbf{F})}^{+})=-\sup_{F\in\mathcal{D}_{n}(\overline{\mathbf{F}})}q_{1-\alpha}^{-}(F)=-\sup_{F\in\mathcal{D}_{n}(\overline{\mathbf{F}})}q_{1-\alpha}^{+}(F)=-H_{\overline{\mathbf{F}}}^{-1}(\alpha),~{}\alpha\in(0,1).$
@@
Note that $H_{\overline{\mathbf{F}}}(x)$ is strictly decreasing and continuous over $(-\infty,\sum_{i=1}^{n}q_{1}^{-1}(\overline{F}_{i}))$ and is equal to 00 on $[\sum_{i=1}^{n}q_{1}^{-1}(\overline{F}_{i}),\infty)$. Hence, we have $F_{\mathcal{D}_{n}(\mathbf{F})}^{+}(x)=H_{\overline{\mathbf{F}}}(-x),~{}x\in\mathbb{R}$.
@@
Next, we show the attainability of $\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $F_{\mathcal{D}_{n}(\mathbf{F})}^{+}$. For any $x\in\mathbb{R}$, there exists a sequence of $F_{k}\in\mathcal{D}_{n}(\mathbf{F}),k\geqslant 1$ such that $\lim_{k\to\infty}F_{k}(x-)=\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)$. For each $F_{k}$, there exists a copula $C_{k}$ such that the joint distribution of $(X_{1},\dots,X_{n})$ is $C_{k}(F_{1},\dots,F_{n})$ and $X_{1}+\dots+X_{n}\sim F_{k}$. It follows from Theorem 25.10 of Billingsley (1995) that we could find a subsequence $\{C_{k_{m}},m\geqslant 1\}$ and a copula $C$ such that $C_{k_{m}}\to C$ weakly as $m\to\infty$, which implies $F_{k_{m}}\to F$ weakly as $m\to\infty$, where $(X_{1},\dots,X_{n})\sim C(F_{1},\dots,F_{n})$ and $X_{1}+\dots+X_{n}\sim F$. Hence, we have $F\in\mathcal{D}_{n}(\mathbf{F})$. Note that there exists a sequence of $y_{l}\uparrow x$ such that $F$ is continuous at all $y_{l}$. Hence, we have $F(y_{l})=\lim_{m\to\infty}F_{k_{m}}(y_{l})\leqslant\liminf_{m\to\infty}F_{k_{m}}(x-)=\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)$. Letting $l\to\infty$, it follows that $F(x-)\leqslant\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)$, implying the attainability of $\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}$. We can similarly show the attainability of $\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{+}$. ∎
@@
Note that the expressions of the extreme probabilities in Proposition 14 are essentially the dual bounds in Theorem 4.17 of Rüschendorf (2013). In light of Theorem 1, Propositions 3 and 10, we obtain the following result.
@@
For $\mathbf{F}\in\mathcal{M}_{D}^{n}\cup\mathcal{M}_{I}^{n}$ and $\Lambda:\mathbb{R}\to[0,1]$, let $F_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $F_{\mathcal{D}_{n}(\mathbf{F})}^{+}$ be given by (14). Then we have
@@
    $\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$; $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(F_{\mathcal{D}_{n}(\mathbf{F})}^{+})$; $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{+})$;
@@
    If $\Lambda$ is decreasing, then (i) remains true by replacing $\tilde{q}_{\Lambda}^{-}$ by $q_{\Lambda}^{-}$, $q_{\Lambda}^{+}$ by $\tilde{q}_{\Lambda}^{+}$ and $q_{\Lambda}^{-}$ by $\tilde{q}_{\Lambda}^{-}$; Moreover, $\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}\tilde{q}_{\Lambda}^{+}(F)=\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$.
@@
is minimized, where $\rho$ is a risk measure and $\left(X_{1},\ldots,X_{n}\right)$ represents the negative returns of $n$ different assets. Here we suppose $X_{1},\dots,X_{n}$ have the identical distribution $F$ as we aim to check whether diversification can reduce the risk.
Recall that a doubly stochastic matrix is a square matrix with nonnegative entries and the sum of each column/row is equal to 1. Let $\mathcal{Q}_{n}$ denote the set of all $n\times n$ doubly stochastic matrices. For two portfolio positions $\boldsymbol{w},\boldsymbol{\gamma}\in\Delta_{n}$, we can say that $\boldsymbol{\gamma}$ is more diversified than $\boldsymbol{w}$, denoted by $\boldsymbol{\gamma}\prec\boldsymbol{w}$, if $\boldsymbol{\gamma}=\Lambda\boldsymbol{w}$ for some $\Lambda\in\mathcal{Q}_{n}$. Note that this binary relationship is also called the majorization order; see e.g., Marshall (2011). Here, we suppose the marginal distribution $X_{i}\sim F$ is known but the dependence structure is completely unknown. Hence, we consider the worst-case scenario. That is to find the optimal portfolio position $\boldsymbol{w}\in\Delta_{n}$ such that
@@
Suppose $\rho=\tilde{q}_{\Lambda}^{+}$ with decreasing $\Lambda$ or $\rho=\tilde{q}_{\Lambda}^{-}$, and $\left(X_{1},\ldots,X_{n}\right)$ has identical marginal distributions $F\in\mathcal{M}_{D}\cup\mathcal{M}_{I}$. If $\boldsymbol{\gamma}\prec\boldsymbol{w}$ , then $\overline{R}_{\rho}(\boldsymbol{\gamma})\geqslant\overline{R}_{\rho}(\boldsymbol{w})$.

### v2

@@ 4.3 Marginal constraints in risk aggregation
The worst-case value of quantiles over the aggregation set has been studied extensively in the literature. It is well known that the worst-case value of quantiles in risk aggregation generally does not admit any analytical formula. Typically, the results in the literature either offer explicit expressions for some special type of marginal distributions or provide some upper bounds for the general marginal distributions. We refer to Embrechts et al. (2006) for bounds, and Wang and Wang (2016), Jakobsons et al. (2016) and Blanchet et al. (2024) for the analytical expressions.
We first offer a variant version of Theorem 1 and Proposition 3, which would be useful to our results in this subsection. For $\Lambda:\mathbb{R}\to[0,1]$, let $\alpha_{\Lambda}^{-}=\inf_{x\in\mathbb{R}}\Lambda(x)$ and $\alpha_{\Lambda}^{+}=\sup_{x\in\mathbb{R}}\Lambda(x)$. The following observation follows directly from the definition of $\Lambda$-quantiles.
Lemma 2.
For an increasing function $f:\mathbb{R}\to[0,1]$ and $\rho\in\{q_{\Lambda}^{-},q_{\Lambda}^{+},\tilde{q}_{\Lambda}^{-},\tilde{q}_{\Lambda}^{+}\}$, if $\beta_{1}<\alpha_{\Lambda}^{-}$ and $\beta_{2}>\alpha_{\Lambda}^{+}$, then we have $\rho(f)=\rho(\beta_{1}\vee f)=\rho(\beta_{2}\wedge f)$.
For $t\in\mathbb{R}$, we say that $t\vee F_{\mathcal{M}}^{-}$ (resp. $t\wedge F_{\mathcal{M}}^{+}$) is attainable if for any $z\in\mathbb{R}$, there exists $F\in\mathcal{M}$ such that $t\vee F(z)=t\vee F_{\mathcal{M}}^{-}(z)$ (resp. $t\wedge F(z)=t\wedge F_{\mathcal{M}}^{+}(z)$). Based on the above lemma, we have the following version of Theorem 1 and Proposition 3.
@@
For $\beta_{1}<\alpha_{\Lambda}^{-}$ and $\beta_{2}>\alpha_{\Lambda}^{+}$, Theorem 1 and Proposition 3 are still valid if $F_{\mathcal{M}}^{-}$ is replaced by $\beta_{1}\vee F_{\mathcal{M}}^{-}$, $F_{\mathcal{M}}^{+}$ by $\beta_{2}\wedge F_{\mathcal{M}}^{+}$ and $\widehat{F}_{\mathcal{M}}^{-}$ by $\beta_{1}\vee\widehat{F}_{\mathcal{M}}^{-}$.
Proof.
By Lemma 2, we immediately obtain that (i) of Theorem 1 still holds when $F_{\mathcal{M}}^{-}$ is replaced by $\beta_{1}\vee F_{\mathcal{M}}^{-}$ and $F_{\mathcal{M}}^{+}$ by $\beta_{2}\wedge F_{\mathcal{M}}^{+}$. Furthermore, using Lemma 2 and the attainability of $\beta_{1}\vee F_{\mathcal{M}}^{-}$ or $\beta_{2}\wedge F_{\mathcal{M}}^{+}$, we can show that (ii) of Theorem 1 still holds after the replacement. The proof is almost the same as before and hence it is omitted. With the replacement in place, (iii) of Theorem 1 follows directly from (i) and (ii) of Theorem 1, along with the fact that $\Lambda$ is decreasing. Finally, the proof of Proposition 3 is almost the same as before and hence, it is omitted. ∎
We say that a distribution $F$ has a monotone density beyond its $t$-quantile for $t\in[0,1)$ if $\frac{(F(x)-t)_{+}}{1-t}$ has a monotone density on its support. Similarly, we say that a distribution $F$ has a monotone density below its $t$-quantile for $t\in(0,1]$ if $\frac{F(x)\wedge t}{t}$ has a monotone density on its support. Let $\mathcal{M}_{t,D}^{+}$ (resp. $\mathcal{M}_{t,I}^{+}$) denote the set of all distributions with decreasing (resp. increasing) densities beyond their corresponding $t$-quantiles, and $\mathcal{M}_{t,D}^{-}$ (resp. $\mathcal{M}_{t,I}^{-}$) denote the set of all distributions with decreasing (resp. increasing) densities below their corresponding $t$-quantiles. The expressions of the extremal distributions over the uncertainty set $\mathcal{D}_{n}(\mathbf{F})$ are displayed in the following proposition.
Proposition 11.
Suppose that $F_{1}^{-1},\dots,F_{n}^{-1}$ are continuous over $(0,1)$. For $\mathbf{F}\in\left(\mathcal{M}_{t,D}^{+}\right)^{n}\cup\left(\mathcal{M}_{t,I}^{+}\right)^{n}$ with $t\in[0,1)$, we have
$t\vee F_{\mathcal{D}_{n}(\mathbf{F})}^{-}=t\vee(1-H_{\mathbf{F}}),$ (14)
@@
with $\mathbf{r}=(r_{1},\dots,r_{n})$, $r=\sum_{i=1}^{n}r_{i}$ and $\Theta_{n}=\{(r_{1},\dots,r_{n})\in\mathbb{R}^{n}:\sum_{i=1}^{n}r_{i}<x\}$.
For $\mathbf{F}\in\left(\mathcal{M}_{t,D}^{-}\right)^{n}\cup\left(\mathcal{M}_{t,I}^{-}\right)^{n}$ with $t\in(0,1]$, we have
$t\wedge F_{\mathcal{D}_{n}(\mathbf{F})}^{+}(x)=t\wedge H_{\overline{\mathbf{F}}},$ (15)
where $\overline{\mathbf{F}}=(\overline{F}_{1},\dots,\overline{F}_{n})$ with $\overline{F}_{i}(x)=\lim_{y\downarrow x}1-F_{i}(-y),~{}x\in\mathbb{R}$.
Moreover, both of $t\vee\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $t\wedge F_{\mathcal{D}_{n}(\mathbf{F})}^{+}$ are continuous and attainable.
@@
First, we suppose that $t\in(0,1)$. In light of Theorems 2 and 4 of Blanchet et al. (2024), we have
$\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{+}(F)=H_{\mathbf{F}}^{-1}(1-\alpha),~{}\alpha\in[t,1),$ (16)
where $H_{\mathbf{F}}^{-1}(\alpha)=\inf\{x\in\mathbb{R}:H_{\mathbf{F}}(x)<\alpha\},~{}\alpha\in(0,1)$. Using the fact $F_{1}^{-1},\dots,F_{n}^{-1}$ are continuous over $(0,1)$, by Lemma 4.5 of Bernard et al. (2014), we have
@@
Moreover, in light of (i) of Corollary 1, we have $\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$ for $\alpha\in(0,1)$. Consequently, $q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})=H_{\mathbf{F}}^{-1}(1-\alpha),~{}\alpha\in[t,1).$ As it is stated in the proof of Theorem 4 of Blanchet et al. (2024), $H_{\mathbf{F}}(x)$ is continuous over $\mathbb{R}$ and strictly decreasing over $(-\infty,\sum_{i=1}^{n}q_{1}^{-1}(F_{i}))$ and is equal to 00 on $[\sum_{i=1}^{n}q_{1}^{-1}(F_{i}),\infty)$. Hence, we have $F_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)=1-H_{\mathbf{F}}(x),~{}x\geqslant H_{\mathbf{F}}^{-1}(1-t)$, which implies that $t\vee F_{\mathcal{D}_{n}(\mathbf{F})}^{-}=t\vee(1-H_{\mathbf{F}})$. We can similarly show that the conclusion holds for $t=0$.
@@
Suppose $t\in(0,1)$. By (i) of Corollary 1, we have $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{+}(F)=q_{\alpha}^{+}(F_{\mathcal{D}_{n}(\mathbf{F})}^{+})$ for $\alpha\in(0,1)$. Direct computation gives $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{+}(F)=-\sup_{F\in\mathcal{D}_{n}(\overline{\mathbf{F}})}q_{1-\alpha}^{-}(F)$, where $\overline{\mathbf{F}}=(\overline{F}_{1},\dots,\overline{F}_{n})$. Note that $\overline{\mathbf{F}}\in\left(\mathcal{M}_{1-t,D}^{-}\right)^{n}\cup\left(\mathcal{M}_{1-t,I}^{-}\right)^{n}$. Hence, using (16) and (17), we have
$q_{\alpha}^{+}(F_{\mathcal{D}_{n}(\mathbf{F})}^{+})=-\sup_{F\in\mathcal{D}_{n}(\overline{\mathbf{F}})}q_{1-\alpha}^{-}(F)=-\sup_{F\in\mathcal{D}_{n}(\overline{\mathbf{F}})}q_{1-\alpha}^{+}(F)=-H_{\overline{\mathbf{F}}}^{-1}(\alpha),~{}\alpha\in(0,t].$
@@
Note that $H_{\overline{\mathbf{F}}}(x)$ is strictly decreasing and continuous over $(-\infty,\sum_{i=1}^{n}q_{1}^{-1}(\overline{F}_{i}))$ and is equal to 00 on $[\sum_{i=1}^{n}q_{1}^{-1}(\overline{F}_{i}),\infty)$. Hence, we have $F_{\mathcal{D}_{n}(\mathbf{F})}^{+}(x)=H_{\overline{\mathbf{F}}}(-x),~{}x\leqslant-H_{\overline{\mathbf{F}}}^{-1}(t)$. This implies $t\wedge F_{\mathcal{D}_{n}(\mathbf{F})}^{+}(x)=t\wedge H_{\overline{\mathbf{F}}}$. We can similarly show that the conclusion holds for $t=1$.
@@
Clearly, $t\vee\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $t\wedge F_{\mathcal{D}_{n}(\mathbf{F})}^{+}$ are continuous. Next, we show the attainability of $\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $F_{\mathcal{D}_{n}(\mathbf{F})}^{+}$, which directly implies the attainability of $t\vee\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $t\wedge F_{\mathcal{D}_{n}(\mathbf{F})}^{+}$. For any $x\in\mathbb{R}$, there exists a sequence of $F_{k}\in\mathcal{D}_{n}(\mathbf{F}),k\geqslant 1$ such that $\lim_{k\to\infty}F_{k}(x-)=\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)$. For each $F_{k}$, there exists a copula $C_{k}$ such that the joint distribution of $(X_{1},\dots,X_{n})$ is $C_{k}(F_{1},\dots,F_{n})$ and $X_{1}+\dots+X_{n}\sim F_{k}$. It follows from Theorem 25.10 of Billingsley (1995) that we can find a subsequence $\{C_{k_{m}},m\geqslant 1\}$ and a copula $C$ such that $C_{k_{m}}\to C$ weakly as $m\to\infty$, which implies $F_{k_{m}}\to F$ weakly as $m\to\infty$, where $(X_{1},\dots,X_{n})\sim C(F_{1},\dots,F_{n})$ and $X_{1}+\dots+X_{n}\sim F$. Hence, we have $F\in\mathcal{D}_{n}(\mathbf{F})$. Note that there exists a sequence of $y_{l}\uparrow x$ such that $F$ is continuous at all $y_{l}$. Hence, we have $F(y_{l})=\lim_{m\to\infty}F_{k_{m}}(y_{l})\leqslant\liminf_{m\to\infty}F_{k_{m}}(x-)=\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)$. Letting $l\to\infty$, it follows that $F(x-)\leqslant\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)$, implying the attainability of $\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{-}$. We can similarly show the attainability of $\widehat{F}_{\mathcal{D}_{n}(\mathbf{F})}^{+}$. ∎
@@
It is worthwhile to mention that the expressions of the extremal distributions in Proposition 11 are essentially the dual bounds in Theorem 4.17 of Rüschendorf (2013). In light of Propositions 10 and 11, we obtain the following result.
@@
Suppose $F_{1}^{-1},\dots,F_{n}^{-1}$ are continuous over $(0,1)$. For $\mathbf{F}\in\left(\mathcal{M}_{t,D}^{+}\right)^{n}\cup\left(\mathcal{M}_{t,I}^{+}\right)^{n}$ with $0\leqslant t<\alpha_{\Lambda}^{-}$, let $t\vee F_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ be given by (14). Then we have
@@
    $\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(t\vee F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$;
@@
    If $\Lambda$ is decreasing, then (i) remains true by replacing $\tilde{q}_{\Lambda}^{-}$ by $q_{\Lambda}^{-}$; Moreover, $\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}\tilde{q}_{\Lambda}^{+}(F)=\sup_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(t\vee F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$.
For $\mathbf{F}\in\left(\mathcal{M}_{t,D}^{-}\right)^{n}\cup\left(\mathcal{M}_{t,I}^{-}\right)^{n}$ with $\alpha_{\Lambda}^{+}<t\leqslant 1$, let $t\wedge F_{\mathcal{D}_{n}(\mathbf{F})}^{+}$ be given by (15). Then we have
  1. (i)
    $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(t\wedge F_{\mathcal{D}_{n}(\mathbf{F})}^{+})$; $\inf_{F\in\mathcal{D}_{n}(\mathbf{F})}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(t\wedge F_{\mathcal{D}_{n}(\mathbf{F})}^{+})$;
  2. (ii)
    If $\Lambda$ is decreasing, then (i) remains true by replacing $q_{\Lambda}^{+}$ by $\tilde{q}_{\Lambda}^{+}$ and $q_{\Lambda}^{-}$ by $\tilde{q}_{\Lambda}^{-}$.
Note that in (i) of Theorem 4, the marginals have monotone densities beyond their corresponding $t$-quantiles for $0\leqslant t<\alpha_{\Lambda}^{-}$. If $\alpha_{\Lambda}^{-}$ is chosen to be sufficiently large such as $\alpha_{\Lambda}^{-}=0.9$ or $0.95$, then we can choose $t$ to be very close to $1$ such as $t=0.89$ or $0.94$. In this case, (i) of Theorem 4 is valid for many commonly used marginal distributions in finance and insurance such as normal, lognormal, $t$, exponential, gamma, and Pareto distributions.
@@
is minimized, where $\rho$ is a risk measure and $\left(X_{1},\ldots,X_{n}\right)$ represents the negative returns of $n$ different assets. Here we suppose that $X_{1},\dots,X_{n}$ have the identical distribution $F$ as we aim to check whether diversification can reduce the risk.
Recall that a doubly stochastic matrix is a square matrix with nonnegative entries and the sum of each column/row is equal to 1. Let $\mathcal{Q}_{n}$ denote the set of all $n\times n$ doubly stochastic matrices. For two portfolio positions $\boldsymbol{w},\boldsymbol{\gamma}\in\Delta_{n}$, we can say that $\boldsymbol{\gamma}$ is more diversified than $\boldsymbol{w}$, denoted by $\boldsymbol{\gamma}\prec\boldsymbol{w}$, if $\boldsymbol{\gamma}={A}\boldsymbol{w}$ for some ${A}\in\mathcal{Q}_{n}$. Note that this binary relationship is also called the majorization order; see e.g., Marshall (2011). Here, we suppose the marginal distribution $X_{i}\sim F$ is known but the dependence structure is completely unknown. Hence, we consider the worst-case scenario. That is to find the optimal portfolio position $\boldsymbol{w}\in\Delta_{n}$ such that
@@
Suppose that $\rho=\tilde{q}_{\Lambda}^{+}$ with decreasing $\Lambda$ or $\rho=\tilde{q}_{\Lambda}^{-}$. Moreover, suppose that $\left(X_{1},\ldots,X_{n}\right)$ has an identical marginal distribution $F\in\mathcal{M}_{t,D}^{+}\cup\mathcal{M}_{t,I}^{+}$ with $0\leqslant t<\alpha_{\Lambda}^{-}$ and $F^{-1}$ is continuous over $(0,1)$. If $\boldsymbol{\gamma}\prec\boldsymbol{w}$ , then $\overline{R}_{\rho}(\boldsymbol{\gamma})\geqslant\overline{R}_{\rho}(\boldsymbol{w})$.